OpenStudy (anonymous):

find the first four terms of the sequence represented by the expression. 3n+5 5, 8, 11, 14 8, 11, 14, 17 3, 6, 9, 12 0, 8, 11, 14

hartnn (hartnn):

to get 1st term, plug in n=1 in 3n+5 to get 2nd term, plug in n=2 in 3n+5 to get 3rd term, plug in n=3 in 3n+5 to get 4th term, plug in n=4 in 3n+5

OpenStudy (anonymous):

8, 11, 14, 17

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

thats correct :)